# Enhancing Agent Learning through World Dynamics Modeling

### v1

@@ 3.2 Online Strategy Learning
To effectively integrate the learned world dynamics $\mathcal{W}$ into the Crafter environments, we deploy an LLM-based agent defined by $\pi:\mathcal{S}\times\mathcal{W}\rightarrow\mathcal{P}(\mathcal{A})$. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ denotes the action space, and $\mathcal{P}$ symbolizes the probability distribution over the action space. Rather than directly mapping the world dynamics $\mathcal{W}$ and the observation $o_{t}$ of the current state $s_{t}$ to the action $a_{t}$, we address the challenge of long-horizon planning by enhancing decision-making with an online strategy learning method. This approach decomposes the planning process into three distinct tasks: sub-goal planning, sub-task planning, and action planning.
@@
Given that the sub-goal sequence $SG=[sg_{1},sg_{2},\ldots]$ is derived from human demonstrations $H$ and treated as a fixed sequence, we employ a simple heuristic for sub-goal planning. When a sub-goal is completed, the current sub-goal will be updated as the first uncompleted sub-goal within $SG$.
@@
For a given current sub-goal $sg_{i}$, we have developed an LLM-based sub-task planner. This planner evaluates and ranks all sub-tasks $st\in ST$ based on the learned world dynamics $\mathcal{W}$, the verbalized current observation $\tilde{o}_{t}$, and the most recently planned sub-task $st_{t-1}$. The top-ranked sub-task is then set as the current sub-task $st_{t}$. To ensure precise execution, the completion of a sub-task $st$ is contingent upon meeting its specific termination condition. This condition is assessed by querying an LLM with the current verbalized observation, the verbalized observation from the time step when the sub-task began, and the termination conditions of the current sub-task.
Online Strategy Deriving
In addition to learning the basic rules of the Crafter environment (i.e., world dynamics $\mathcal{W}$), we also explore the development of advanced game-playing strategies based on these dynamics. Unlike learning the world dynamics, the strategy space is often too vast to fully explore. To address this, we propose evolving the dynamics into strategies $\mathcal{I}$ using an online learning method. This method reduces the search space by conditioning not only on the dynamics $\mathcal{W}$ but also on the verbalized current observation $\tilde{o}_{t}$ and the sub-task $st_{i}$. This targeted approach helps generate strategies that are more grounded and responsive to current game scenarios than those developed through offline methods. To facilitate this, we have designed an LLM-based Evolver that develops strategy candidates $\tilde{\mathcal{I}}$ through deductive reasoning applied to $\mathcal{W}$. Specifically, the Evolver derives strategy candidates by rules of inference, such as modus ponens. These strategy candidates $\tilde{\mathcal{I}}$ are evaluated for their validity and are ranked based on their usefulness by an LLM-based critic. Lastly, the valid and useful candidates are added to the situational strategy set $\mathcal{I}$.
@@
The final action selection process is carried out in two primary steps:
@@
    Invalid Action Masking: This step involves masking all actions that are not feasible under the current game conditions, based on the verified dynamics $\mathcal{W}$ and current observation $\tilde{o}_{t}$.
@@
    Action Selection: From the set of valid actions, a specific primitive action $a$ is chosen. This selection is based on several factors: the current sub-task $st_{i}$, the verbalized current observation $\tilde{o}_{t}$, the world dynamics $\mathcal{W}$, a windowed history of previously planned actions and observations, and the strategies derived $\mathcal{I}$.
4 Experiment
In this section, we start by detailing our experimental settings including evaluation metrics and the set of baselines we are comparing against (Section 4.1). Then, we report the main results (Section 4.2). Additionally, we conduct a set of controlled experiments, we provide quantitative and qualitative analyses to obtain better understanding on DiVE (Section 4.3 and 4.4).
4.1 Environment Settings
@@
Environment Crafter Hafner (2021) is a set of open-world survival games on $64\times 64$ grid maps, featuring a diverse array of materials such as tree, stone, and coal, as well as entities including cow, zombie, and skeleton semi-randomly spawn on the maps. The games include an achievement graph with 22 unique achievements across 7 levels. The agent observes its surroundings through a local $7\times 9$ window and is also aware of its status within the game environment.
Verbalizer
@@
The text description from the verbalizer includes: the nearest object of each type within the accumulated observations, the objects located between these nearest objects, the objects in each direction, and the agent’s current inventory and status. We provide an example of the verbalization process in Appendix D.
Evaluation metrics
@@
Following previous works Hafner (2021); Wu et al. (2024b), we evaluate the agent with two primary metrics: reward and score. Agents receive a $+1$ reward for each new achievement unlocked (e.g., make wood pickaxe, place furnace). Additionally, they earn a ±$\pm 0.1$ reward for every health point gained or lost. The score metric is computed by aggregating the success rates for each achievement:
$S\doteq\exp\left(\frac{1}{N}\sum_{i=1}^{N}\ln\left(1+s_{i}\right)\right)-1.$ (4)
Baseline systems
We compare DiVE with three categories of baselines. First, LLM-based methods, including SPRING Wu et al. (2024b), ELLM Du et al. (2023), and chain-of-thought (CoT) Wei et al. (2022). Second, reinforcement learning (RL) based approaches such as DreamerV3 Hafner et al. (2023), PPO Schulman et al. (2017), AD  Moon et al. (2024). Third, we consider demonstrations from human experts, which provide insights from actual human players.
@@
Table 1: Primary table illustrating the performance of DVE in comparison to all baseline models. The symbol indicates results obtained by running the official code implementation using the same five random seed and same model.
4.2 Overall Results
Table 1 demonstrates that DiVE significantly outperforms all other baselines across two distinct evaluation metrics. Notably, DiVE exceeds the previous state-of-the-art (SOTA) LLM-based method, SPRING [^1] Results were obtained by executing the official SPRING code using the same LLM configuration (GPT-4o) and environment random seeds as DiVE, ensuring a fair comparison., by a substantial margin, achieving a 337.8% relative improvement in score and a 110.1% enhancement in reward. Additionally, DiVE also surpasses the prior SOTA RL-based approach, DreamerV3, with a 21.4% absolute improvement in score and a 2.8 absolute increase in reward. Notably, DiVE achieves rewards comparable to domain-familiar human players using just 10 human demonstrations.
4.3 Quantitative Analysis
@@
Table 2: Different components’ impact on the performance; The abbreviations D, V, and E stand for Discover, Verifier, and Evolver, respectively; S means the dynamics (manual) used in the SPRING paper; H means the human-annotated dynamics.
Ablation study
@@
We conduct a set of ablation study to help elucidate how individual elements contribute to the overall performance of the method. As shown in the first section of Table 2, various variants of DiVE have been designed to assess the effectiveness of each component within the methodology. Notably, the significant performance disparity between DiVE and its variant lacking the Evolver component empirically underscores the Evolver’s role in effectively developing gameplay strategies based on world dynamics $\mathcal{W}$, thereby enhancing the agent’s proficiency in the Crafter environment. Similarly, the performance decline observed in the variant without the Verifier highlights the importance of formulating strategies $\mathcal{I}$ based on accurate world dynamics $\mathcal{W}$. Moreover, the additional performance reduction in the version missing both Verifier and Evolver demonstrates their orthogonal functions—precision in dynamics from the Verifier and advancement in strategies from the Evolver.
@@
The performance of DiVE without Discoverer, Verifier, and Evolver reverts to the CoT baseline, indicating that merely decomposing the task according to the hierarchy $\mathcal{H}$ without incorporating domain knowledge $\mathcal{K}_{\text{target}}$ offers no advantage. The significant gap between CoT + D&V and DiVE w/o E further evinces that an LLM-based agent struggles in long-horizon planning tasks without task decomposition, emphasizing the necessity of the decomposition hierarchy $\mathcal{H}$.
Sources of dynamics
@@
We also explore how DiVE performs with world dynamics derived from various sources. As shown in the second section of Table 2, DiVE significantly outperforms the variants that utilize dynamics $S^{{\dagger}}$ as used in the SPRING paper Wu et al. (2024b). This performance improvement suggests that the learned dynamics $\mathcal{W}$ are more beneficial than $S^{{\dagger}}$, possibly because $S^{{\dagger}}$ lacks some of the helpful details present in $\mathcal{W}$. We will delve deeper into this aspect in upcoming case studies. The performance difference between methods using $S^{{\dagger}}$ with and without the Evolver also underscores the importance of strategy evolution, the effectiveness of which is closely linked to the quality of the world dynamics it employs.
@@
Besides dynamics learned from human demonstrations and the manual dynamics used in SPRING, we have also investigated a third source: human-annotated dynamics. The findings reveal that DiVE performs comparably to the variants that use human-annotated dynamics, demonstrating the robustness and effectiveness of DiVE’s approach to dynamic learning.
@@
Figure 3: Recall of learned dynamics over discovery steps with mean and standard deviation.
Quality of learned dynamics
As previously mentioned, directly quantifying the desired properties is challenging since we cannot precisely measure how much domain-relevant information $\mathcal{K}_{\text{rel}}$ an LLM contains or determine the exact amount of knowledge $\mathcal{K}_{\text{target}}$ required by a domain. However, by using human-annotated dynamics $H^{{\dagger}}$ as a silver standard for $\mathcal{K}_{\text{target}}$ , we can estimate the precision and recall of the learned dynamics $\mathcal{W}$. This approach allows us to effectively track and evaluate the progress of the LLM-based dynamic learning. Specifically, we calculate recall as $R=\frac{|\mathcal{W}\cap H^{{\dagger}}|}{|H^{{\dagger}}|}$ and precision as $P=\frac{|\mathcal{W}\cap H^{{\dagger}}|}{|\mathcal{W}|}$. As illustrated in Figure 3, both the discovered dynamics $\tilde{\mathcal{W}}$ and the verified dynamics $\mathcal{W}$ demonstrate an increase in recall with the progression of discovery steps, indicating that the richness of the learned dynamics enhances as more discovery steps are taken. Furthermore, the gap in recall between $\tilde{\mathcal{W}}$ and $\mathcal{W}$ narrows as the number of discovery steps increases, suggesting that the Verifier effectively filters out "noisy" dynamic candidates while retaining those that generalize across different trajectory segments.
@@
To further investigate whether the Verifier retains correct world dynamic candidates $\tilde{\mathcal{W}}{+}$ while filtering out the inaccurate ones $\tilde{\mathcal{W}}{-}$, we examine the precision of both the discovered dynamics $\tilde{\mathcal{W}}$ and the verified dynamics $\mathcal{W}$. As shown in Figure 4, the precision of the verified dynamics consistently and significantly surpasses that of the discovered dynamics, demonstrating that the Verifier effectively identifies and eliminates inaccurate dynamic candidates. This validates the Verifier’s role in enhancing the reliability of the dynamics used for decision-making processes.
Figure 4: Precision of learned dynamics before and after dynamic verification.
4.4 Qualitative Analysis

### v2

@@ 3.2 Online Strategy Learning
To effectively incorporate the learned world dynamics $\mathcal{W}$ into the downstream domains, we deploy an LLM-based agent defined by $\pi:\mathcal{S}\times\mathcal{W}\rightarrow\mathcal{P}(\mathcal{A})$. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ denotes the action space, and $\mathcal{P}$ symbolizes the probability distribution over the action space. Instead of directly mapping the world dynamics $\mathcal{W}$ and the current state observation $o_{t}$ to the action $a_{t}$, we tackle the challenge of long-horizon planning by integrating an online strategy learning method. This approach decomposes the planning process into three distinct stages: sub-goal planning, sub-task planning, and action planning.
@@
Given that the sub-goal sequence $SG=[sg_{1},sg_{2},\ldots]$ is derived from human demonstrations $H$ and treated as a fixed sequence, we utilize a straightforward heuristic for sub-goal planning. When a sub-goal is completed, the current sub-goal is updated to the first uncompleted sub-goal in $SG$.
@@
For a given current sub-goal $sg_{i}$, we have developed an LLM-based sub-task planner. This planner evaluates and ranks all sub-tasks $st\in ST$ based on the learned world dynamics $\mathcal{W}$, the verbalized current observation $\tilde{o}_{t}$, and the most recently executed sub-task $st_{t-1}$. The highest-ranked sub-task is then designated as the current sub-task $st_{t}$. To ensure accurate execution, the completion of a sub-task $st$ is contingent upon satisfying its specific termination condition. This condition is verified by querying an LLM using the current verbalized observation, the observation at the time the sub-task began, and the termination conditions of the current sub-task.
Learning Strategies
In addition to learning the fundamental rules of the downstream domains, we also focus on developing advanced game-playing strategies based on these dynamics. Unlike world dynamics learning, the strategy space is often too expansive for exhaustive exploration. To address this challenge, we propose evolving the dynamics into strategies, denoted as $\mathcal{I}$, using an online learning approach.
This method reduces the search space by conditioning not only on the dynamics $\mathcal{W}$, but also on the verbalized current observation $\tilde{o}_{t}$ and the sub-task $st_{i}$. This targeted approach enables the generation of strategies that are more contextually grounded and responsive to current game scenarios compared to those developed through offline methods. To support this process, we have designed an LLM-based Evolver that generates strategy candidates $\tilde{\mathcal{I}}$ by applying deductive reasoning to the learned dynamics $\mathcal{W}$. Specifically, the Evolver derives strategy candidates using rules of inference, such as modus ponens. These strategy candidates, denoted as $\tilde{\mathcal{I}}$, are then evaluated for validity and ranked based on their utility by an LLM-based critic. Finally, the valid and useful candidates are incorporated into the situational strategy set $\mathcal{I}$.
@@
The final action selection process is executed in two main steps:
@@
    Invalid Action Masking: This step involves filtering out actions that are infeasible under the current situation, based on the verified dynamics $\mathcal{W}$ and current verbalized observation $\tilde{o}_{t}$.
@@
    Action Selection: From the set of valid actions, a specific primitive action $a$ is chosen based on multiple factors: the current sub-task $st_{i}$, the verbalized current observation $\tilde{o}_{t}$, the world dynamics $\mathcal{W}$, a windowed history of previously planned actions and observations, and the derived strategies $\mathcal{I}$.
4 Experiment Setup
To demonstrate the effectiveness of DiVE in bridging the knowledge gap, we evaluate its performance within the Crafter and MiniHack environments. For a fair comparison, all LLM-based agents, including ours, utilize the GPT-4o model and the same environment seed in the Crafter setting. Further details on the setups can be found in Appendix A.1 and B.1.
4.1 Crafter
@@
Crafter Hafner (2021) is an open-world survival game set on $64\times 64$ grid-based maps, featuring a diverse array of materials such as tree, stone, and coal, as well as entities including cow, zombie, and skeleton semi-randomly spawn on the maps. The games include an achievement graph with 22 unique achievements across 7 levels. The agent perceives its surroundings through a local $7\times 9$ observation window and maintains awareness of its status within the game environment.
@@
The text description generated by the verbalizer includes: the nearest object of each type within the accumulated observations, the objects situated between these nearest objects, the objects in each direction, as well as the agent’s current inventory and status. An example of the verbalization process is provided in Appendix B.5.
@@
The agent is evaluated using two primary metrics: reward and score. Agents receive a $+1$ reward for each new achievement unlocked (e.g., make wood pickaxe, place furnace) and a ±$\pm 0.1$ reward for every health point gained or lost. The score is calculated by aggregating the success rates $s_{i}$ for each achievement and is formulated as:
$S\doteq\exp\left(\frac{1}{N}\sum_{i=1}^{N}\ln\left(1+s_{i}\right)\right)-1.$
We compare DiVE against the following baselines:
  • LLM-based approaches: SPRING Wu et al. (2024b), ELLM Du et al. (2023), and Chain-of-Thought (CoT)Wei et al. (2022).
  • Reinforcement Learning (RL) approaches: DreamerV3Hafner et al. (2023), PPO Schulman et al. (2017), and AD Moon et al. (2024).
  • Human Players: Expert performance on the Crafter environment.
4.2 MiniHack
MiniHack Samvelyan et al. (2021) is a grid-based environment built on the video game NetHack Küttler et al. (2020). Unlike the Crafter environment, it supports the creation of tasks that target specific agent capabilities. In our work, we focus on the Skill Acquisition Tasks subset, which evaluates the agent’s ability to leverage the rich diversity of NetHack’s objects, monsters, and dungeon features, as well as their interactions. These tasks introduce complex world dynamics, where actions are factorized autoregressively and require executing a sequence of follow-up actions for the initial action to produce the desired effect.
We use the Lava Crossing, Wand of Death, and Quest tasks as testbeds to evaluate DiVE. We provide detailed descriptions of these task in Appendix A.2. The agent is rewarded for unlocking achievements within each task. To facilitate interaction with LLMs, we use the NLE language wrapper Goodger et al. (2023) to verbalize both observations and actions from the environment.
We compare DiVE against SSO Nottingham et al. (2024) and Reflexion Shinn et al. (2024), both of which require prior training on the tasks. Following the official guidelines, we run each method for 30 iterations, evaluating performance after every 10 iterations by attempting the task 10 times using a fixed set of skills or reflections. The final performance is determined based on the results from the last 10 evaluation attempts.
5 Experimental Results
@@
Table 1: Performance comparison of DiVE against baseline models in the Crafter environment. Methods with indicate that they were obtained using the official code implementation, executed with the same five random seeds and model configurations as used for DiVE.
We evaluate the performance of DiVE in the Crafter and MiniHack environments. In Section 5.1, we present the overall results to demonstrate DiVE’s effectiveness in bridging the knowledge gap in these tasks. Section 5.2 provides a detailed analysis of the contributions of individual components through controlled experiments. Then, in Section 5.3, we evaluate the effectiveness of the dynamics learned by DiVE in the Crafter environment. Finally, in Section 5.4, we further analyze the learned dynamics in both the Crafter environment and a modified MiniHack setting with altered dynamics.
5.1 DiVE’s Performance
Table 1 and  2 demonstrate that DiVE surpasses all other baselines in the Crafter and MiniHack environments. In the Crafter environment, DiVE exceeds the previous state-of-the-art (SOTA) LLM-based method, SPRING, by a substantial margin, achieving a 337.8% relative improvement in score and a 110.1% enhancement in reward. Additionally, DiVE also surpasses the prior SOTA RL-based approach, DreamerV3, with a 21.4% absolute improvement in score and a 2.8 absolute increase in reward. Notably, DiVE achieves rewards comparable to human players using 10 demonstrations.
In the MiniHack environment, with only a single demonstration, DiVE matches the performance of SSO and Reflexion (both of which require 30 iterations of training) on the Lava Crossing task, and outperforms both baselines on the Wand of Death and Quest tasks. Specifically, DiVE achieves a 68% improvement over Reflexion on the Wand of Death task and a 30% and 36% improvement over SSO and Reflexion, respectively, on the Quest task.
Method Lava Crossing Wand of Death Quest
Reflexion 0.8$\pm$0.0 0.32$\pm$0.15 0.87$\pm$0.43
SSO 0.8$\pm$0.0 0.52$\pm$0.31 0.91$\pm$0.29
DiVE 0.8$\pm$0.0 0.54$\pm$0.34 1.18$\pm$0.38
Table 2: Comparison of DiVE performance against baseline models in the MiniHack environment. Results marked with were obtained using the official code implementation and evaluated over 10 runs.
5.2 Contribution of Individual Components
@@
Table 3: Impact of different components on performance: Crafter. D, V, and E represent Discover, Verifier, and Evolver, respectively; S refers to dynamics derived from the game manual, H refers to human-annotated dynamics.
We conduct a series of ablation studies to clarify the contribution of each individual element to DiVE’s overall performance. We report the results on Crafter in the first section of Table 3, and the results on MiniHack in Table 4.
Crafter
@@
The significant performance gap between DiVE and its variant without the Evolver component empirically demonstrates the Evolver’s effectiveness in developing gameplay strategies based on world dynamics $\mathcal{W}$, thereby enhancing the agent’s overall proficiency in this environment. Similarly, the performance decline observed in the variant without the Verifier underscores the importance of formulating strategies $\mathcal{I}$ based on accurate world dynamics $\mathcal{W}$. Moreover, the further performance drop in the version lacking both the Verifier and Evolver components highlights their complementary roles—the Verifier ensures precision in capturing dynamics, while the Evolver focuses on strategy development.
@@
The performance of DiVE without the Discoverer, Verifier, and Evolver components reverts to the CoT baseline, indicating that simply decomposing the task according to the hierarchy $\mathcal{H}$ without integrating domain knowledge $\mathcal{K}_{\text{target}}$ provides no performance benefit. The substantial gap between CoT + D&V and DiVE w/o E further demonstrates that an LLM-based agent struggles with long-horizon planning tasks in the absence of task decomposition, underscoring the importance of the decomposition hierarchy $\mathcal{H}$.
MiniHack
Since MiniHack is built on the popular video game NetHack Goodger et al. (2023), LLMs already possess a certain level of understanding of the environment. For example, DiVE without the Discover, Verifier, and Evolver components (CoT baseline) can successfully solve the Lava Crossing task. However, as the complexity of the tasks increases, using LLMs alone is insufficient for solving more challenging scenarios like the Wand of Death and Quest tasks, highlighting a potential knowledge gap for these more intricate problems.
The primary performance gains of DiVE in the MiniHack environment can be attributed to the Discover component, which accurately identifies the underlying environment dynamics required for each task. Due to the embedded knowledge of LLMs about the MiniHack environment, they can identify these dynamics effectively, as shown by the minimal performance drop when DiVE is used without the Verifier. However, in the modified MiniHack setting described in Section 5.4, we highlight the critical role of the Verifier in maintaining reliability. Additionally, since the tasks have a relatively short horizon, the dynamics uncovered by the Discover component are sufficient to complete them successfully.
Components Lava Crossing Wand of Death Quest
DiVE 0.80$\pm$0.00 0.54$\pm$0.34 1.18$\pm$0.38
w/o E 0.72$\pm$0.24 0.52$\pm$0.30 1.16$\pm$0.48
w/o V 0.69$\pm$0.25 0.52$\pm$0.38 1.18$\pm$0.38
w/o V&E 0.72$\pm$0.24 0.40$\pm$0.40 1.11$\pm$0.48
w/o D&V&E 0.72$\pm$0.17 0.20$\pm$0.00 0.41$\pm$0.29
Table 4: Impact of different components on performance: MiniHack. D, V, and E represent Discover, Verifier, and Evolver, respectively.
5.3 Evaluation of Learned Dynamics
@@
We investigate the performance of DiVE leveraging world dynamics derived from different sources in Crafter. As shown in the second section of Table 3, DiVE significantly outperforms variants that utilize the dynamics $S^{{\dagger}}$ from the game manual Wu et al. (2024b). This performance improvement indicates that the learned dynamics $\mathcal{W}$ are more advantageous than $S^{{\dagger}}$, likely because $S^{{\dagger}}$ lacks certain beneficial details that are captured in $\mathcal{W}$. The performance gap between methods using $S^{{\dagger}}$ with and without the Evolver further highlights the importance of strategy evolution, whose effectiveness is closely tied to the quality of the underlying world dynamics.
@@
In addition to dynamics learned from human demonstrations and game manual, we have explored a third source: human-annotated dynamics. The results show that DiVE performs comparably to the variants using human-annotated dynamics, demonstrating the robustness and effectiveness of DiVE’s approach to dynamic learning.
As previously mentioned, it is difficult to quantify the desired properties because we cannot precisely measure the domain-relevant information $\mathcal{K}_{\text{relevant}}$ in LLMs or the exact amount of domain knowledge $\mathcal{K}_{\text{target}}$ required. However, by using human-annotated dynamics $H^{{\dagger}}$ as a reference benchmark for $\mathcal{K}_{\text{target}}$, we can estimate the precision and recall of the learned dynamics $\mathcal{W}$, enabling us to effectively assess the progress of LLM-based dynamic learning. Specifically, we define recall as $R=\frac{|\mathcal{W}\cap H^{{\dagger}}|}{|H^{{\dagger}}|}$ and precision as $P=\frac{|\mathcal{W}\cap H^{{\dagger}}|}{|\mathcal{W}|}$.
@@
Figure 3: Recall of learned dynamics over discovery steps, presented with mean and standard deviation, in the Crafter environment.
As illustrated in Figure 3, both the discovered dynamics $\tilde{\mathcal{W}}$ and the verified dynamics $\mathcal{W}$ exhibit an increase in recall as the number of discovery steps progresses, indicating that the richness of the learned dynamics improves over time. Moreover, the narrowing gap in recall between $\tilde{\mathcal{W}}$ and $\mathcal{W}$ suggests that the Verifier effectively filters out ’noisy’ dynamic candidates while preserving those that generalize across different trajectory segments.
@@
To evaluate whether the Verifier preserves correct world dynamic candidates $\tilde{\mathcal{W}}{+}$ while filtering out unreliable ones $\tilde{\mathcal{W}}{-}$, we analyze the precision of both the discovered dynamics $\tilde{\mathcal{W}}$ and the verified dynamics $\mathcal{W}$. As shown in Figure 4, the precision of the verified dynamics consistently and significantly exceeds that of the discovered dynamics, demonstrating the Verifier’s effectiveness in identifying and eliminating inaccurate candidates. This confirms the Verifier’s role in enhancing the reliability of the dynamics used for decision-making.
5.4 Analysis of Learned Dynamics